# Learning in Wilson-Cowan model for metapopulation

### v1

@@ Abstract
The Wilson-Cowan model for metapopulation, a Neural Mass Network Model, treats different subcortical regions of the brain as connected nodes, with connections representing various types of structural, functional, or effective neuronal connectivity between these regions. Each region comprises interacting populations of excitatory and inhibitory cells, consistent with the standard Wilson-Cowan model. By incorporating stable attractors into such a metapopulation model’s dynamics, we transform it into a learning algorithm capable of achieving high image and text classification accuracy. We test it on MNIST and Fashion MNIST, in combination with convolutional neural networks, on CIFAR-10 and TF-FLOWERS, and, in combination with a transformer architecture (BERT), on IMDB, always showing high classification accuracy. These numerical evaluations illustrate that minimal modifications to the Wilson-Cowan model for metapopulation can reveal unique and previously unobserved dynamics.
@@
[2]organization=Department of Mathematics and Computer Science, University of Perugia,addressline=Via Vanvitelli 1, city=Perugia, postcode=06123, state=Perugia, country=Italy
@@ 1 Introduction
Understanding brain information processing requires building computational models that are capable of performing cognitive tasks (Kriegeskorte and Douglas, 2018). A brain computational model is a mathematical model that mimics the brain information processing underlying the performance of some task at some level of abstraction. At microscale, biological processes that underlie brain computation are described by biophysical models (Murray et al., 2018; Pathak et al., 2022), while, at macroscale, processes occurring in the brain are modeled by brain-dynamical and causal-interaction models (Karameh and Nahas, 2019; Sun et al., 2014). In the past decades, the field of computational neuroscience has developed many mathematical models of elementary computational components (McCulloch and Pitts, 1943; Rosenblatt, 1958) and their implementation with biological neurons (Dayan and Abbott, 2005; Abbott, 2008). In this paper, we discuss one of them: the Wilson-Cowan model (Wilson and Cowan, 1972).
@@
The Wilson-Cowan model describes the evolution of excitatory and inhibitory activity in a synaptically coupled neuronal network. As opposed to being a detailed biophysical model, it is a coarse-grained description of the overall activity of a large-scale neuronal network, employing just two differential equations (Kilpatrick, 2022). As such, they embraced nonlinear dynamics, but in an interpretable form, i.e., they were motivated by physiological evidence (Mountcastle, 1997), which suggested the existence of certain populations of neurons with similar responses to external stimuli (Hubel and Wiesel, 1965). Indeed, it was the first mathematical formulation to emphasize the significance of interactions between excitatory (E) and inhibitory (I) neural populations in cortical tissue (Harms and Elias, 2013), thereby incorporating both cooperation and competition (Wilson and Cowan, 2021). A core part of the Wilson–Cowan modeling framework is the use of a sigmoid function to determine population firing rates in terms of population activity. Although the use of a sigmoid is now ubiquitous throughout computational neuroscience, in the original Wilson–Cowan model formulation this was assumed to arise via some form of averaging over heterogeneity (or noise) in networks of simple threshold elements (Byrne et al., 2020).
Before the Wilson-Cowan model, the nascent field of computational neuroscience was still grappling with how best to model neural dynamics. Thereafter, it began to coalesce onto a common set of neural network equations, and studying their properties became the focus (Chow and Karimipanah, 2020). Moreover, the model introduced many ideas from dynamical systems (Strogatz, 2018) to the wider community, highlighting that switching, cycling, and information storage could all be viewed using the framework of attractor dynamics (Ashwin et al., 2024; Yu et al., 2024; Zhang et al., 2024; Khona and Fiete, 2022).
@@
This model has been widely used to study various aspects of neural activity: as a single-node description of excitatory-inhibitory population dynamics, as a building block for larger-scale brain network modeling studies, and as the underpinning of spatially extended models of neural dynamics at the tissue scale. It takes into account essential parameters such as the strength of synaptic connections among each type of neuronal population and the intensity of input received by each population. By manipulating these parameters, the model can replicate a range of dynamic brain behaviors, including multistability (Kaslik et al., 2024) and limit cycles (Wilson and Cowan, 1972). A sufficient condition for the existence of multiple steady states, for instance, is that the average number of synapses between excitatory neurons must exceed the variance of the distribution of these connections (Wilson and Cowan, 1972). Conversely, the existence of limit cycles in a neural population requires that the interactions within the excitatory subpopulation be significantly stronger than those within the inhibitory subpopulation (Wilson and Cowan, 1972). Other applications are: stable inhomogeneous steady states that store information dynamically and suggest a basis for short-term memory (Byrne, 2017), oscillations (Maruyama et al., 2014), traveling waves (Harris and Ermentrout, 2018), and the formation of spatial patterns (Wilson and Cowan, 1973). Additionally, the model captures information processing (Campbell and Wang, 1996), binocular rivalry (Wilson, 2017), cognitive dynamics of movement (Erlhagen and Schöner, 2002), phase-amplitude coupling (Duchet et al., 2020), neuroimaging data (González-Ramírez, 2024; Corbitt et al., 2018), and cortical resonant frequencies (Cowan et al., 2016; Lea-Carnall et al., 2016), among other complex brain activities (Painchaud et al., 2022; Destexhe and Sejnowski, 2009; Kora and Simon, 2024).
For example, the seminal work of Ermentrout and Cowan (1979), considering a Wilson–Cowan model in two spatial dimensions (representing the surface of visual cortex), explains the emergence of visual hallucination patterns in terms of a Turing instability (Turing, 2004). More recent studies, introducing multiple excitatory populations, demonstrated the presence of long range excitatory–excitatory interconnections among neurons with multiple orientation preferences located at a distance from each other and aligned roughly collinearly that permits a much wider range of visual hallucinations to be explained (Bressloff et al., 2001).
The Wilson–Cowan equations have also been applied to epilepsy and decision. For epilepsy, this was based upon the original observations that spatially localized limit cycles could exist and that travelling waves could occur should the inhibition be too weak (Wilson and Cowan, 1973; Shusterman and Troy, 2008; Meijer et al., 2015). The results were comparable to data recorded from cortical surface electrodes during passage of epileptic seizures. For decision, instead, a candidate network, presumably mimicking areas in the prefrontal lobe (Binder et al., 2009), has been proposed related to the Wilson–Cowan network for rivalry (Wilson, 2013). Lastly, Wilson–Cowan models are a key component of the Virtual Brain project that aims to deliver the first simulation of the human brain based on personalized large-scale connectivity (Sanz-Leon et al., 2015).
@@
Building upon these foundational methods, the Wilson–Cowan model has been extended providing a deeper insight into the emergent collective behaviors of networks across multiple scales of organization (Sarracino et al., 2020; Nandi et al., 2023; de Candia et al., 2021). Most obviously, it is possible to generalize to multiple excitatory and inhibitory populations reflective of particular cortical areas and functions. For examples, Conti and Van Gorder (2019) studied the dynamics of a network of Wilson–Cowan model (a system of connected Wilson–Cowan oscillators). By observing that information transfer within each cortical area is not instantaneous, they consider a system of delay differential equations with two different kinds of discrete delay for exploring a variety of larger networks, in order to determine how the network topology will influence time delayed Wilson–Cowan dynamics. They find that network structure can regularize or deregularize the dynamics. In Sanchez-Rodriguez et al. (2024), the authors, instead, use personalized Alzheimer’s disease computational models built on whole-brain Wilson-Cowan oscillators to evaluate the direct impact of toxic protein deposition on neuronal activity.
@@
To fill this gap, in this paper, we consider a network metapopulation form of the Wilson-Cowan model (also known as a Neural Mass Network Model (Byrne et al., 2020)). Such a network treats different subcortical regions of the brain as connected nodes, and the connections represent various types of links between them, as structural, functional, or effective connectivity between these distinct subcortical regions (DeWeerdt, 2019). Within each region, or node of the network, there are interacting subpopulations of excitatory and inhibitory cells (Bazinet et al., 2023), in accord with the standard Wilson-Cowan model (a cartoon is displayed in Fig. 1). Noting that since long-range connections in the brain mainly project from excitatory pyramidal cells (Bekkers, 2011; Gerfen et al., 2018), we restrict couplings between brain regions to connections between only excitatory subpopulations. By planting stable attractors into the dynamics of this model (Marino et al., 2023), we aim to explore the learning capabilities and generalization abilities of such a network on classification tasks. Specifically, our approach leverages machine learning techniques to learn the coupling matrix of our Wilson-Cowan model for metapopulation, enabling us to validate classification performance of such a model (and combinations with Convolutional Neural Networks (CNNs)) on MNIST dataset (Deng, 2012) (grey scale handwritten images of size $28\times 28$), Fashion-MNIST dataset (Xiao et al., 2017) (grey scale Zalando’s article images of size $28\times 28$), CIFAR-10 dataset (Krizhevsky et al., 2009) (colored images of size $32\times 32\times 3$) and TF-FLOWERS dataset (TensorFlow-Team, 2019) (colored images of size $224\times 224\times 3$). Moreover, we tested our model in combination with a transformer architecture for classifying text reviews from the IMDB dataset (Pal et al., 2020). Although this method, also in combination with other computational neural models, does not reach the state-of-the-art deep learning algorithms for classification tasks, it illustrates that minimal modifications to the Wilson-Cowan model for metapopulation can unveil unique and previously unobserved dynamics. Our research, therefore, aims to shed light on how the presence of planted attractors within this Neural Mass Network Model affects its ability to learn and generalize on different classification tasks.
@@ 2 The model
According to graph theory, brain networks can be described as graphs composed of nodes (vertices) representing neural elements (neurons or brain regions) linked by edges representing physical connections (synapses or axonal projections).
The topological and physical distances between elements in brain networks are intricate. Neurons and brain regions, containing populations of neurons, have a higher probability of being connected if they are spatially close, whereas connections between spatially remote neurons or brain regions are less likely (Bullmore and Sporns, 2009). This is because longer axonal projections are more expensive in terms of material and energy costs; indeed, the spatial layout of neurons or brain regions is economically arranged to minimize axonal volume (Bullmore and Sporns, 2012).
@@
The interactions between vertices are described, according to graph theory terminology, by an adjacency matrix $\mathbf{A}$ of the graph underlying the network. In general $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$. In this manuscript, an element of a matrix $\mathbf{A}$ is defined as $[\mathbf{A}]_{ln}$. The matrix $\mathbf{A}$, in our case, can be factorized by definition, i.e., $\mathbf{A}=\Phi\Lambda\Phi^{-1}$, where $\Lambda$ is a diagonal matrix $\Lambda\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$, composed by the eigenvalues of $\mathbf{A}$, and $\Phi\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is a matrix where each column, i.e., $\vec{\phi}_{n}=[\Phi]_{:,n}$, with $n=1,\dots,\mathcal{N}$, is an eigenvector of $\mathbf{A}$. With the symbol $(\cdot)^{-1}$ we define the inverse matrix operation. The choice of dealing with the above decomposition of the coupling matrix echoes the spectral approach to Deep Learning discussed in Chicchi et al. (2024) and reference therein.
@@
We define $\gamma$ as the ratio between inhibitors and activators in the population of neurons on a single vertex of the graph. This ratio is conventionally set to $0.25$. This is due to the fact that it is known that in the brain, there are about $50$ billion neurons in the cortex. Among them, scientists have shown that $80\%$ are excitatory, whereas the remaining $20\%$ are inhibitory (Markram et al., 2004), depending on whether they have a prominent postsynaptic density or a very thin postsynaptic density, respectively (DeFelipe and Fariñas, 1992).
@@
where $\vec{e}\in\mathbb{R}^{\mathcal{N}}$ is a vector where all components are set to $1$, $\Gamma=1/\sqrt{\mathcal{N}}$ is a scaling factor (Hopfield, 1982; Marino and Ricci-Tersenghi, 2024; Huang, 2021), $\omega_{EE}$, $\omega_{II}$, $\omega_{EI}$, and $\omega_{IE}$ denote the strength of connectivity within and between the excitatory and inhibitory subpopulations. $h_{E}$ and $h_{I}$ denote the strength of the external input to each of these subpopulations. The $\Gamma\mathbf{A}\vec{z}_{1}(t)$ term couples the nodes of the network, and gives the strength of the interactions between connected nodes (Pietras and Daffertshofer, 2019). We recall that $\mathbf{A}\vec{z}_{1}(t)$ is a multiplication between a matrix in $\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and a vector in $\mathbb{R}^{\mathcal{N}}$. It is worth pointing out, as well described in the Introduction, that inter-node connections are established between excitatory subpopulations only, as is true of the brain. Such a term has been introduced in many works (Butler et al., 2012; Abeysuriya et al., 2018; Ahmadizadeh et al., 2016; Pinder and Crofts, 2021) for analyzing different dynamics in Wilson-Cowan oscillator networks, including approaches that use a Laplacian instead of an adjacency matrix. In all these cases, the authors analyze only the dynamical behaviors, without embedding stable attractors or applying the resultant dynamics to a learning task. As far as we know, our work is the first in this direction of research.
The system’s evolution can lead to unique fixed points, multiple fixed points each with its own basin of attraction, or even result in an oscillating periodic attractor, depending on the choice of parameters in equations (1) and (2). Throughout this paper, our primary focus will be on parameter configurations leading to bistability (Alvankar Golpayegan and de Candia, 2023) (Fig. 2 displays a cartoon of the bistable profile for the standard Wilson-Cowan model). We select the bistabile profile to construct a dictionary for our classification task. Specifically, we aim to utilize the fixed points of the oscillator as targets for the dynamics to converge to. We, therefore, select parameters in equation (1) and (2) to ensure the emergence of precisely two fixed points at every node of the graph (in the case the coupling term is omitted, i.e., when $\Gamma=0$), as used in Zankoc et al. (2017). We denote these solutions as $(\overline{x}^{{}^{(1)}},\overline{y}^{{}^{(1)}})$ and $(\overline{x}^{{}^{(2)}},\overline{y}^{{}^{(2)}})$ for each node.
@@
In the framework of a bistable system, with the matrix $\mathbf{A}$ being factorizable ( $\mathbf{A}=\Phi\Lambda\Phi^{-1}$), we embed attractors into the system dynamics, as presented in Marino et al. (2023).
Our objective is to classify $K$ items based on the output of the network $\mathcal{G}(\mathcal{N},\mathcal{E})$. Having in mind that, we encode $K$ attractors within the matrix $\Phi$, representing distinct system configurations that it may evolve towards over time. We specify $K$ columns of $\Phi$, thus $\vec{\phi}_{k}$ for $k=1,\dots,K$. By design, these column vectors are chosen permutations of the fixed points identified in the dynamics described by equation (1), such that each element of $\vec{\phi}_{k}$ corresponds to either $\overline{x}^{{}^{(1)}}$ or $\overline{x}^{{}^{(2)}}$, namely $[\vec{\phi}_{k}]_{i}=$$\{\overline{x}^{{}^{(1)}}$ or $\overline{x}^{{}^{(2)}}\}$ $\forall i=1,\dots,\mathcal{N}$ (see Fig. 4, Panel b) and c)). Furthermore, we align only the eigenvectors $\vec{\phi}_{k}$, $k=1,\dots,K$, within the kernel of $\mathbf{A}$, highlighting that these vectors, as elements of the kernel, can be constructed from linear combinations of the kernel’s orthonormal vectors.
In conclusion, $\mathbf{A}\vec{\phi}_{k}=\lambda_{k}\vec{\phi}_{k}=\vec{0}$. Thus, each eigenvalue $\lambda_{k}$ associated to a $\vec{\phi}_{k}$ ($k=1,\dots,K$) in $\Lambda$ is set to zero. To streamline our approach without sacrificing generality, we choose the system parameters so that $\overline{y}=\overline{y}^{{}^{(1)}}=\overline{y}^{{}^{(2)}}$. Although $\overline{y}$ represents a degree of freedom within the system dynamics, it lacks informational value for our classification task. Consequently, we refrain from adding an eigenvector composed exclusively of $\overline{y}$ values to $\mathbf{A}$.
The embedded eigenvectors must be stable. To ensure this, we perform a linear stability analysis (Strogatz, 2018) to identify the conditions that the non-zero eigenvalues in $\Lambda$ must satisfy in order to make $\vec{\phi}_{k}$ ($k=1,\dots,K$) stable attractors.

### v2

@@ Abstract
The Wilson-Cowan model for metapopulation, a Neural Mass Network Model, treats different subcortical regions of the brain as connected nodes, with connections representing various types of structural, functional, or effective neuronal connectivity between these regions. Each region comprises interacting populations of excitatory and inhibitory cells, consistent with the standard Wilson-Cowan model. In this paper, we show how to incorporate stable attractors into such a metapopulation model’s dynamics. By doing so, we transform the Neural Mass Network Model into a biologically inspired learning algorithm capable of solving different classification tasks. We test it on MNIST and Fashion MNIST in combination with convolutional neural networks, as well as on CIFAR-10 and TF-FLOWERS, and in combination with a transformer architecture (BERT) on IMDB, consistently achieving high classification accuracy.
@@
[2]organization=Department of Mathematics and Computer Science, University of Florence, addressline= Viale Morgagni 67, city=Firenze, postcode=50134, state=Firenze, country=Italy
@@ 1 Introduction
Understanding brain information processing requires building computational models that are capable of performing cognitive tasks [50]. A brain computational model is a mathematical model that mimics the brain information processing underlying the performance of some task at some level of abstraction. At microscale, biological processes that underlie brain computation are described by biophysical models [70, 75], while, at macroscale, processes occurring in the brain are modeled by brain-dynamical and causal-interaction models [43, 89]. In the past decades, the field of computational neuroscience has developed many mathematical models of elementary computational components [64, 81] and their implementation with biological neurons [1]. In this paper, we discuss one of them: the Wilson-Cowan model [94].
@@
The Wilson-Cowan model describes the evolution of excitatory and inhibitory activity in a synaptically coupled neuronal network. As opposed to being a detailed biophysical model, it is a coarse-grained description of the overall activity of a large-scale neuronal network, employing just two differential equations [45]. As such, they embraced nonlinear dynamics, but in an interpretable form, i.e., they were motivated by physiological evidence [69], which suggested the existence of certain populations of neurons with similar responses to external stimuli [42]. Indeed, it was the first mathematical formulation to emphasize the significance of interactions between excitatory (E) and inhibitory (I) neural populations in cortical tissue [36], thereby incorporating both cooperation and competition [96].
@@
This model has been widely used to study various aspects of neural activity: as a single-node description of excitatory-inhibitory population dynamics, as a building block for larger-scale brain network modeling studies, and as the underpinning of spatially extended models of neural dynamics at the tissue scale. It takes into account essential parameters such as the strength of synaptic connections among each type of neuronal population and the intensity of input received by each population. By manipulating these parameters, the model can replicate a range of dynamic brain behaviors, including multistability [44] and limit cycles [94]. Other applications are: stable inhomogeneous steady states that store information dynamically and suggest a basis for short-term memory [15], oscillations [63], traveling waves [37], and the formation of spatial patterns [95]. Additionally, the model captures information processing [16], binocular rivalry [93], cognitive dynamics of movement [29], phase-amplitude coupling [28], neuroimaging data [35], cortical resonant frequencies [19], epilepsy [95, 65], and decision [9], among other complex brain activities [72, 48]. Lastly, Wilson–Cowan models are a key component of the Virtual Brain project that aims to deliver the first simulation of the human brain based on personalized large-scale connectivity [83].
@@
Building upon these foundational methods, the Wilson–Cowan model has been extended providing a deeper insight into the emergent collective behaviors of networks across multiple scales of organization [71]. Most obviously, it is possible to generalize to multiple excitatory and inhibitory populations reflective of particular cortical areas and functions. For examples, [18] studied the dynamics of a network of Wilson–Cowan model (a system of connected Wilson–Cowan oscillators). By observing that information transfer within each cortical area is not instantaneous, they consider a system of delay differential equations with two different kinds of discrete delay for exploring a variety of larger networks, in order to determine how the network topology will influence time delayed Wilson–Cowan dynamics. They find that network structure can regularize or deregularize the dynamics. In [82], the authors, instead, use personalized Alzheimer’s disease computational models built on whole-brain Wilson-Cowan oscillators to evaluate the direct impact of toxic protein deposition on neuronal activity.
@@
To fill this gap, in this paper, we consider a network metapopulation form of the Wilson-Cowan model (a Neural Mass Network Model [14]). Biologically speaking, such a network treats different subcortical regions of the brain as connected nodes, and the connections represent various types of links between them, as structural, functional, or effective connectivity between these distinct subcortical regions [26]. Within each region, or node of the network, there are interacting subpopulations of excitatory and inhibitory cells [6], in accord with the standard Wilson-Cowan model (a cartoon is displayed in Fig. 1). Moreover, noting that since long-range connections in the brain mainly project from excitatory pyramidal cells [7, 33], we restrict couplings between brain regions to connections between only excitatory subpopulations.
Given such a network, we aim to explore the learning capabilities and generalization abilities on classification tasks. To do so, we present, as in [58], a method for planting stable attractors into the dynamics of this Neural Mass Network Model (as described in Sec. 2, 3, 4 and illustrated in Fig. 4). The presented method incorporates stable attractors into the model’s dynamics, identifiable as targets for the classification task. These attractors are pairs of eigenvectors and eigenvalues of the adjacency matrix of the graph, selected a priori to meet the task requirements. Each eigenvector, used as long-term memory, consists of stable fixed points of the dynamics of each node. When the model reaches an attractor, the dynamics stabilizes, positioning on one of the targets of the classification task.
Using machine learning techniques to learn the others eigenvectors and eigenvalues of the matrix, we validate the classification performance of such a model (and combinations with Convolutional Neural Networks (CNNs)) on MNIST dataset [23] (grey scale handwritten images of size $28\times 28$), Fashion-MNIST dataset [97] (grey scale Zalando’s article images of size $28\times 28$), CIFAR-10 dataset [51] (colored images of size $32\times 32\times 3$) and TF-FLOWERS dataset [91] (colored images of size $224\times 224\times 3$). Moreover, we tested our model in combination with a transformer architecture for classifying text reviews from the IMDB dataset [73].
@@ 2 The model
According to graph theory, brain networks can be described as graphs composed of nodes (vertices) representing neural elements (brain regions) linked by edges representing physical connections (synapses or axonal projections).
The topological and physical distances between elements in brain networks are intricate. Neurons and brain regions, containing populations of neurons, have a higher probability of being connected if they are spatially close, whereas connections between spatially remote neurons or brain regions are less likely [11]. This is because longer axonal projections are more expensive in terms of material and energy costs; indeed, the spatial layout of neurons or brain regions is economically arranged to minimize axonal volume [12].
@@
The interactions between vertices are described, according to graph theory terminology, by an adjacency matrix $\mathbf{A}$ of the graph underlying the network. In general $\mathbf{A}\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$. In this manuscript, an element of a matrix $\mathbf{A}$ is defined as $[\mathbf{A}]_{ln}$. The matrix $\mathbf{A}$, in our case, can be factorized by definition, i.e., $\mathbf{A}=\Phi\Lambda\Phi^{-1}$, where $\Lambda$ is a diagonal matrix $\Lambda\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$, composed by the eigenvalues of $\mathbf{A}$, and $\Phi\in\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ is a matrix where each column, i.e., $\vec{\phi}_{n}=[\Phi]_{:,n}$, with $n=1,\dots,\mathcal{N}$, is an eigenvector of $\mathbf{A}$. With the symbol $(\cdot)^{-1}$ we define the inverse matrix operation. The choice of dealing with the above decomposition of the coupling matrix echoes the spectral approach to Deep Learning discussed in [17] and reference therein.
@@
We define $\gamma$ as the ratio between inhibitors and activators in the population of neurons on a single vertex of the graph. This ratio is conventionally set to $0.25$. This is due to the fact that it is known that in the brain, there are billions of neurons in the cortex. Among them, scientists have shown that $80\%$ are excitatory, whereas the remaining $20\%$ are inhibitory [62], depending on whether they have a prominent postsynaptic density or a very thin postsynaptic density, respectively [21].
@@
where $\vec{e}\in\mathbb{R}^{\mathcal{N}}$ is a vector where all components are set to $1$, $\Gamma=1/\sqrt{\mathcal{N}}$ is a scaling factor [61, 40], $\omega_{EE}$, $\omega_{II}$, $\omega_{EI}$, and $\omega_{IE}$ denote the strength of connectivity within and between the excitatory and inhibitory subpopulations. $h_{E}$ and $h_{I}$ denote the strength of the external input to each of these subpopulations. The $\Gamma\mathbf{A}\vec{z}_{1}(t)$ term couples the nodes of the network, and gives the strength of the interactions between connected nodes [76, 57]. We recall that $\mathbf{A}\vec{z}_{1}(t)$ is a multiplication between a matrix in $\mathbb{R}^{\mathcal{N}\times\mathcal{N}}$ and a vector in $\mathbb{R}^{\mathcal{N}}$. It is worth pointing out that inter-node connections are established between excitatory subpopulations only, as is true of the brain. Such a term has been introduced in many works, as [2, 77], for analyzing different dynamics in Wilson-Cowan oscillator networks, including approaches that use a Laplacian instead of an adjacency matrix. In all these cases, the authors analyze only the dynamical behaviors, without embedding stable attractors or applying the resultant dynamics to a learning task. As far as we know, our work is the first in this direction of research.
The system’s evolution can lead to unique fixed points, multiple fixed points each with its own basin of attraction, or even result in an oscillating periodic attractor, depending on the choice of parameters in equations (1) and (2). Throughout this paper, our primary focus will be on parameter configurations leading to bistability [4] (Fig. 2 displays a cartoon of the bistable profile for the standard Wilson-Cowan model). We select the bistabile profile to construct a dictionary for our classification task. Specifically, we aim to utilize the fixed points of the oscillator as targets for the dynamics to converge to. We, therefore, select parameters in equation (1) and (2) to ensure the emergence of precisely two fixed points at every node of the graph (in the case the coupling term is omitted, i.e., when $\Gamma=0$), as used in [99]. We denote these solutions as $\vec{\overline{\zeta}}^{(1)}=(\overline{x}^{{}^{(1)}},\overline{y}^{{}^{(1)}})^{\intercal}$ and $\vec{\overline{\zeta}}^{(2)}=(\overline{x}^{{}^{(2)}},\overline{y}^{{}^{(2)}})^{\intercal}$ for each node.
@@
Our objective is to classify $K$ items based on the output of the network $\mathcal{G}(\mathcal{N},\mathcal{E})$. The network which defines the backbone of the examined dynamical model is made of $\mathcal{N}$ nodes, $\mathcal{N}$ denoting, for example, the number of pixels of the images to be classified. Each node is assigned with a standard Wilson-Cowan model. Nodes are then sensing its nearest neighbours, as stipulated by a linear coupling term which selectively acts on the excitatory species. The web of inter-nodes connections is stored in the $\mathcal{N}\times\mathcal{N}$ adjacency weighted matrix ($\mathbf{A}=\Phi\Lambda\Phi^{-1}$), whose elements represent the weights to be optimized in the training process. The above matrix is formulated in spectral domain [58]: a subset of $K$ suitably tailored eigenvectors is assigned to its kernel and define the attractive poles for the globally coupled dynamics. The eigenvalues populate a specific region of the real axis that yields stability of the crafted attractors, as dictated by the linear stability analysis presented in the next section. In essence the model to be trained is an extended collection of interacting units subject to contrasting tendencies: local reactions occur between subpopulations referred to a single brain region (or node), while global, spatially (across nodes) extended, interactions make the system to evolve towards complex heterogeneous stable states. In summary, the item to be classified is supplied as an initial condition of the dynamical model and the ensuing stable attractor, as reached by the trained model, flags for the corresponding classification label. Under this perspective, learning amounts to shaping the basin of attraction of the underlying dynamical model.
Having in mind that, we encode $K$ attractors within the matrix $\Phi$, representing distinct system configurations that it may evolve towards over time. We specify $K$ columns of $\Phi$, thus $\vec{\phi}_{k}$ for $k=1,\dots,K$. By design, these column vectors are chosen permutations of the fixed points identified in the dynamics described by equation (1), such that each element of $\vec{\phi}_{k}$ corresponds to either $[\vec{\overline{\zeta}}^{(1)}]_{1}$ or $[\vec{\overline{\zeta}}^{(2)}]_{1}$, namely $[\vec{\phi}_{k}]_{i}=$$\{[\vec{\overline{\zeta}}^{(1)}]_{1}$ or $[\vec{\overline{\zeta}}^{(2)}]_{1}\}$ $\forall i=1,\dots,\mathcal{N}$ (see Fig. 4, Panel b) and c)). Furthermore, we align only the eigenvectors $\vec{\phi}_{k}$, $k=1,\dots,K$, within the kernel of $\mathbf{A}$, highlighting that these vectors, as elements of the kernel, can be constructed from linear combinations of the kernel’s orthonormal vectors.
In conclusion, $\mathbf{A}\vec{\phi}_{k}=\lambda_{k}\vec{\phi}_{k}=\vec{0}$. Thus, each eigenvalue $\lambda_{k}$ associated to a $\vec{\phi}_{k}$ ($k=1,\dots,K$) in $\Lambda$ is set to zero. To streamline our approach without sacrificing generality, we choose the system parameters so that $[\vec{\overline{\zeta}}^{(1)}]_{2}=[\vec{\overline{\zeta}}^{(2)}]_{2}$. Although $[\vec{\overline{\zeta}}^{(1)}]_{2}$ represents a degree of freedom within the system dynamics, it lacks informational value for our classification task. Consequently, we refrain from adding an eigenvector composed exclusively of $[\vec{\overline{\zeta}}^{(2)}]_{2}$ values to $\mathbf{A}$.
The embedded eigenvectors must be stable. To ensure this, we perform a linear stability analysis [88] to identify the conditions that the non-zero eigenvalues in $\Lambda$ must satisfy in order to make $\vec{\phi}_{k}$ ($k=1,\dots,K$) stable attractors.